# Advancing GPU IPC for Stiff Affine-Deformable Simulation

### v2

@@ 7.2. Stress Tests
In this example, we observe significantly faster PCG convergence with a diagonal preconditioner than with the MAS preconditioner in GIPC. Thus, we apply the diagonal preconditioner for both GIPC and ’Ours (SRBK)’. This difference arises because Morton code-based reordering may spread nodes from the same box into different groups. Without connections between boxes, nodes from different boxes in the same group cannot be effectively merged, slowing down MAS performance. In contrast, our CEMAS minimizes inter-group connections, effectively grouping nodes from the same box, enabling faster PCG convergence as shown in ’Ours (CEMAS+SRBK)’. Interestingly, approximating the boxes as stiff affine bodies does not significantly improve performance in this scenario, as each box’s small DOF count allows efficient computation with pure FEM.
@@ 8. Conclusion and Discussion
In this paper, we presented a novel, fully GPU-optimized Incremental Potential Contact (IPC) framework capable of simulating materials of a wide range of stiffness with consistent high performance and scalability. Our framework introduces several significant advancements over previous methods, achieving up to 10$\times$ speedup compared to the state-of-the-art GIPC. Our contributions include: 1) A novel connectivity-enhanced MAS preconditioner optimized for the GPU, which significantly improves convergence at a lower preconditioning cost; 2) A cubic strain-limiting energy with an analytic eigensystem, designed to simulate stiff membranes such as cloth without membrane locking; 3) An innovative hash-based multi-layer reduction strategy, enabling fast matrix assembly for efficient affine-deformable coupling. Our framework’s robust performance is demonstrated through extensive benchmark studies and rigorous performance analyses, showing superior results in various scenarios, including soft, stiff, and hybrid simulations. The framework efficiently handles high resolution, large deformations, and high-speed impacts, maintaining accuracy and robustness. We believe our advancements address critical challenges in IPC simulation, providing a scalable and high-performance solution for complex frictional contact behaviors. Our comprehensive approach not only improves the efficiency of IPC but also retains the method’s reliability and accuracy, making it a valuable tool for future research and applications in a variety of fields.
@@
Our preconditioner was designed under the assumption that the mesh connectivity would not change during the simulation. We precompute the partition when loading the meshes, which means our current design is not suitable for simulations with adaptive mesh refinement and/or fracture mechanics, where the mesh topology may change in any frame. Additionally, the main performance bottleneck of IPC comes from two main factors: the expensive computational cost of each Newton iteration and the large number of Newton iterations required in cases with challenging contact scenarios. Our preconditioner design and GPU optimizations have primarily focused on minimizing the cost of each Newton iteration. However, the slow convergence of the optimization in challenging cases can still result in slower overall performance compared to simulation methods that trade accuracy for efficiency. Therefore, it is very meaningful to keep investigating better preconditioners that can effectively handle simulations with mesh topology changes and to explore novel nonlinear solvers or discretization strategies that can significantly improve the convergence of the nonlinear solver.

### v3

@@ 7.2. Stress Tests
In this example, we observe significantly faster PCG convergence with a diagonal preconditioner than with the MAS preconditioner in GIPC. Thus, we apply the diagonal preconditioner for both GIPC and ‘Ours (SRBK)’. This difference arises because Morton code-based reordering may spread nodes from the same box into different groups. Without connections between boxes, nodes from different boxes in the same group cannot be effectively merged, slowing down MAS performance. In contrast, our CEMAS minimizes inter-group connections, effectively grouping nodes from the same box, enabling faster PCG convergence as shown in ‘Ours (CEMAS+SRBK)’. Interestingly, approximating the boxes as stiff affine bodies does not significantly improve performance in this scenario, as each box’s small DOF count allows efficient computation with pure FEM.
@@ 8. Conclusion and Discussion
In this paper, we presented a novel, fully GPU-optimized Incremental Potential Contact (IPC) framework capable of simulating materials of a wide range of stiffness with consistent high performance and scalability. Our framework introduces several significant advancements over previous methods, achieving up to 10$\times$ speedup compared to the state-of-the-art GIPC. Our contributions include: 1) A novel connectivity-enhanced MAS preconditioner optimized for the GPU, which significantly improves convergence at a lower preconditioning cost; 2) A cubic inexact strain-limiting energy with an analytic eigensystem, designed to simulate stiff membranes such as cloth without membrane locking; 3) An innovative hash-based two-level reduction strategy, enabling fast matrix assembly for efficient affine-deformable coupling. Our framework’s robust performance is demonstrated through extensive benchmark studies and rigorous performance analyses, showing superior results in various scenarios, including soft, stiff, and hybrid simulations. The framework efficiently handles high resolution, large deformations, and high-speed impacts, maintaining accuracy and robustness. We believe our advancements address critical challenges in IPC simulation, providing a scalable and high-performance solution for complex frictional contact behaviors. Our comprehensive approach not only improves the efficiency of IPC but also retains the method’s reliability and accuracy, making it a valuable tool for future research and applications in a variety of fields.
@@
Our GPU optimizations leverage efficient mutual data access among threads within the same warp, making NVIDIA CUDA a natural demonstration platform due to its widespread adoption. These optimizations are generalizable to any platform with warp-like operations; however, without this capability, they would not be feasible. Our preconditioner was designed under the assumption that the mesh connectivity would not change during the simulation. We precompute the partition when loading the meshes, which means our current design is not suitable for simulations with adaptive mesh refinement and/or fracture mechanics, where the mesh topology may change in any frame. Additionally, the main performance bottleneck of IPC comes from two main factors: the expensive computational cost of each Newton iteration and the large number of Newton iterations required in cases with challenging contact scenarios. Our preconditioner design and GPU optimizations have primarily focused on minimizing the cost of each Newton iteration. However, the slow convergence of the optimization in challenging cases can still result in slower overall performance compared to simulation methods that trade accuracy for efficiency. Therefore, it is very meaningful to keep investigating better preconditioners that can effectively handle simulations with mesh topology changes and to explore novel nonlinear solvers or discretization strategies that can significantly improve the convergence of the nonlinear solver.